rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#3695 — FS#7635 — Dep 35 Collect

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
Many customers escalated us browsing problems.
The synchronisation is good but the browsing is difficult. google ok
The sites after the search no.


It seems to be an MTU problem.

We did the tests and tried many LNS. The problem persists.
We suspect a problem on the collect network.
All the customers are on the same LAC.

The investigations are in progress.
Date:  Sunday, 16 December 2012, 19:33PM
Reason for closing:  Done